OpenStudy (anonymous):

If x is a binomial random variable, compute p(x) for each of the following cases: n=5, x=1, p = 0.2 n=4, x=2, q = 0.4 n=3, x=0, p = 0.7 n=5, x=3, p = 0.1 n=4, x= 2, q = 0.6 n=3, x=1, q = 0.9

OpenStudy (anonymous):

P(x)=n!/ (x!(n−x)!) * p^x(1−p)^(n−x) is the formula you have to evaluate for each number set. Can you see where that takes you? the n! Just means 1*2*....*n So plug in the numbers and see what you get.

OpenStudy (anonymous):

Let me walk you through the first example then n=5, x=1, p = 0.2 We know P(x)=n!/ (x!(n−x)!) * p^x(1−p)^(n−x) so 5!/(1! * (5-1)!) * 0.2^1 * (1-0.20)^(5-1) = 120/24 * 0.2 * 0.8^4 = 0.4096